OpenStudy (anonymous):

the measures of the angles of a triangle are in the extended ratio 2:2:8 what is the measure of the smallest angel

OpenStudy (jdoe0001):

\(\large \begin{array}{llllll} 2&:&2&:&8\\ \uparrow &&\uparrow &&\uparrow \\ 2&&2&&2\cdot 4 \end{array}\qquad \begin{array}{llll} \textit{2 angles are identical}\\ \textit{the other is 4 times either of the other 2} \end{array}\) 

OpenStudy (jdoe0001):

hmm meant to be 4x rather ... ok... so \(\large \begin{array}{llllll} 2&:&2&:&8\\ \uparrow &&\uparrow &&\uparrow \\ 2&&2&&2\cdot 4 \end{array}\qquad \begin{array}{llll} \textit{2 angles are identical}\\ \textit{the other is 4 times either of the other 2} \end{array}\)